a waitress has 3 time as many dimes as nickles, one more quarter than nickles that worth $2.65 in total. find how many dimes she has?
Found 2 solutions by lynnlo, stanbon:
Answer by lynnlo(4176)
Answer by stanbon(75887)

a waitress has 3 time as many dimes as nickles, one more quarter than nickles that worth $2.65 in total. find how many dimes she has?
-----------------
Equations:
d = 3n
q = n+1
5n + 10d + 25q = 265 cents
---------------------------------
Substitute for d and for q: solve for "n":
5n + 10(3n) + 25(n+1) = 265
-----
5n + 30n + 25n = 240
==================
60n = 240
n = 6 (# of nickels)
---
d = 3n = 18 (# of dimes)
==============================
Cheers,
Stan H.
================